OpenStudy (anonymous):

y varies inversely as twice x. When x = 2, y = 3. Find y when x = 6

OpenStudy (whpalmer4):

If y varies inversely with 2x, then \[y = \frac{k}{2x}\]where \(k\) is the constant of variation. Plug in your known values to determine the value of \(k\), then use the updated equation to find y when x = 6

OpenStudy (anonymous):

i dont know where to even start.

OpenStudy (whpalmer4):

Take my equation. Plug in the values in the "when " statement. Solve for k. Replace k in my equation with the number you got. Plug in the value of x from the "find y when" statement and, well, find y! :-)

OpenStudy (whpalmer4):

do you understand how/why I wrote that equation?

OpenStudy (anonymous):

ya hold up im working the problem out

OpenStudy (whpalmer4):

np, I'll get a notification when you respond...

OpenStudy (anonymous):

ya i cant figure it out..

OpenStudy (whpalmer4):

okay, we know x=2, y = 3 \[y = \frac{k}{2x}\]\[3=\frac{k}{2*2}\]Solve that for \(k\)

OpenStudy (whpalmer4):

k is the number that when divided by 4 gives you 3. what is k?

OpenStudy (whpalmer4):

dude, are you like growing some extra fingers to work it out, or what?

OpenStudy (anonymous):

6?

OpenStudy (whpalmer4):

6 divided by 4 gives you 3?

OpenStudy (anonymous):

no buyt my options is 1, 2, 4, or 6

OpenStudy (whpalmer4):

we aren't done yet! we're trying to find k, which you need to find the answer.

OpenStudy (whpalmer4):

you shouldn't need to make any reference to the answer key to work this problem out with 100% certainty.

OpenStudy (whpalmer4):

so, I ask again, what number, divided by 4, equals 3?

OpenStudy (anonymous):

12

OpenStudy (whpalmer4):

right! now, we know that k=12, so our equation becomes \[y = \frac{12}{2x}\]All we have to do to get the answer for the problem is plug in the value of x for which they want to know the value of y.

OpenStudy (anonymous):

y=6?

OpenStudy (anonymous):

so x=1?

OpenStudy (whpalmer4):

Yes, x = 1. Here's a picture:

OpenStudy (whpalmer4):

You can see that when x=2, y = 3, and when x = 6, y = 1 (y, not x)

OpenStudy (whpalmer4):

and that's the curve shape you get with inverse variation

OpenStudy (anonymous):

so the final answer is 1

OpenStudy (whpalmer4):

yes.

OpenStudy (anonymous):

thank you.

OpenStudy (whpalmer4):

you're welcome.